Joe rides his bike 10 minutes to his friends house then leaves his bike and walks 15 minutes to the Library. Joe rides his bike 10km/h faster than he walks. If the enter trip is 2.75km how far is it from his friends house to the Library?
Answer by lwsshak3(11628)

Joe rides his bike 10 minutes to his friends house then leaves his bike and walks 15 minutes to the Library. Joe rides his bike 10km/h faster than he walks. If the enter trip is 2.75km how far is it from his friends house to the Library?
**
let x=Joe's walking speed
x+10=Joe's bike speed
distance=travel time*speed
10 min=10/60 hr=1/6 hr
15 min=15/60 hr=1/4 hr
bike distance+walking distance=2.75 km
(1/6)(x+10)+(1/4)(x)=2.75
LCD:12
2x+20+3x=33
5x=13
x=13/5 km/hr (Joe's walking speed
ans:
distance from his friends house to the Library
=speed*travel time=13/5*1/4=.65km